# DeepEye-SQL: A Software-Engineering-Inspired Text-to-SQL Framework

### v1

@@ 1. Introduction
Text-to-SQL is a task that converts natural-language questions into SQL queries over a database (Luo et al., 2025; Liu et al., 2025a; Baek et al., 2025). Large language models (LLMs) (Wu et al., 2025) have substantially advanced Text-to-SQL, achieving strong results on benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2024a). For example, Alpha-SQL (Li et al., 2025b) leverages dynamic multi-step reasoning, while XiYan-SQL (Liu et al., 2025c) improves SQL generation and multi-candidate SQL selection through task-specific fine-tuning.
@@ 2.2. Text-to-SQL Solutions
The task of translating natural language into executable SQL queries has evolved significantly (Liu et al., 2025a). The recent emergence of Large Language Models (LLMs) has marked a new paradigm. Current research on applying LLMs to Text-to-SQL is largely divided into two categories: fine-tuning and prompting-based methods. Fine-tuning methods adapt open-source models like Code Llama or Qwen for the Text-to-SQL task by training them on large corpora of question-SQL pairs (Li et al., 2024c; Yang et al., 2024; Li et al., 2025a; Sheng et al., 2025; Liu et al., 2025c). This approach can yield highly efficient and specialized models but may exhibit limited generalization to out-of-domain or highly complex scenarios. In contrast, prompting-based methods leverage the powerful in-context learning and reasoning capabilities of very large, often proprietary, models like GPT-4o (Hurst et al., 2024) and Gemini (Anil et al., 2023) without requiring model training. To manage the complexity of the task, these methods typically decompose the process into a multi-stage pipeline, including sub-tasks like schema linking, SQL generation, and refinement (Pourreza and Rafiei, 2023; Li et al., 2024b; Pourreza et al., 2025; Li et al., 2025b). While powerful, these frameworks often rely on a single generation path and the fallible self-correction abilities of the LLM, which can compromise robustness. Our work, DeepEye-SQL, belongs to the prompting-based category but differentiates itself by explicitly adopting a systematic framework inspired by software engineering.
@@ 4.4. SQL Unit Testing and Revision via Tool-Chain
Syntax Checker SQL syntax and execution errors SELECT * FORM table (typo)
WHERE column = (incomplete)
JOIN Checker Non-standard JOIN conditions ON cond1 OR cond2
ON column IN (...)
ORDER-BY Checker Invalid ORDER BY + LIMIT combinations ORDER BY MAX(...) LIMIT 1
ORDER BY COUNT(*) LIMIT 3
Time Checker Incorrect time function usage and formatting STRFTIME(’%Y’, date) format
Invalid date comparisons
SELECT Checker Ambiguous return columns SELECT * → specific columns
Unnecessary wildcards
MaxMin Checker Suboptimal MAX/MIN patterns WHERE col = (SELECT MAX(...))
ORDER BY col DESC LIMIT 1
NULL Checker NULL values in ORDER BY columns Add WHERE col IS NOT NULL
for ORDER BY columns
Result Checker Empty or meaningless results Queries returning only NULL
Zero-row result sets
@@ 6. Conclusion
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v2

@@ 1. Introduction
Text-to-SQL is a task that converts natural-language questions into SQL queries over a database (Luo et al., 2025b; Liu et al., 2025a; Luo et al., 2018, 2022, 2021, 2025a; Baek et al., 2025). Large language models (LLMs) (Wu et al., 2025; Shuai et al., 2025; Zhu et al., 2024; Zhang et al., 2025d) have substantially advanced Text-to-SQL, achieving strong results on benchmarks such as Spider (Yu et al., 2018) and BIRD (Li et al., 2024a). For example, Alpha-SQL (Li et al., 2025b) leverages dynamic multi-step reasoning, while XiYan-SQL (Liu et al., 2025c) improves SQL generation and multi-candidate SQL selection through task-specific fine-tuning.
@@ 2.2. Text-to-SQL Solutions
The task of translating natural language into executable SQL queries has evolved significantly (Liu et al., 2025a). The recent emergence of Large Language Models (LLMs) has marked a new paradigm. Current research on applying LLMs to Text-to-SQL is largely divided into two categories: fine-tuning and prompting-based methods. Fine-tuning methods adapt open-source models like Code Llama or Qwen for the Text-to-SQL task by training them on large corpora of question-SQL pairs (Li et al., 2024c; Yang et al., 2024; Li et al., 2025a; Sheng et al., 2025; Liu et al., 2025c). This approach can yield highly efficient and specialized models but may exhibit limited generalization to out-of-domain or highly complex scenarios. In contrast, prompting-based methods leverage the powerful in-context learning and reasoning capabilities of very large, often proprietary, models like GPT-4o (Hurst et al., 2024) and Gemini (Anil et al., 2023) without requiring model training. To manage the complexity of the task, these methods typically decompose the process into a multi-stage pipeline, including sub-tasks like schema linking, SQL generation, and refinement (Pourreza and Rafiei, 2023; Li et al., 2024b; Pourreza et al., 2025; Li et al., 2025b; Ma et al., 2024; Zhu et al., 2025). While powerful, these frameworks often rely on a single generation path and the fallible self-correction abilities of the LLM, which can compromise robustness. Our work, DeepEye-SQL, belongs to the prompting-based category but differentiates itself by explicitly adopting a systematic framework inspired by software engineering.
@@ 4.4. SQL Unit Testing and Revision via Tool-Chain
Syntax Checker SQL syntax and SELECT * FORM table (typo)
execution errors WHERE column = (incomplete)
JOIN Checker Non-standard ON cond1 OR cond2
JOIN conditions ON column IN (...)
ORDER-BY Checker Invalid ORDER BY ORDER BY MAX(...) LIMIT 1
+ LIMIT combinations ORDER BY COUNT(*) LIMIT 3
Time Checker Incorrect time function STRFTIME(’%Y’, date) format
usage and formatting Invalid date comparisons
SELECT Checker Ambiguous return SELECT * → specific columns
columns Unnecessary wildcards
MaxMin Checker Suboptimal MAX/MIN WHERE col = (SELECT MAX(...))
patterns ORDER BY col DESC LIMIT 1
NULL Checker NULL values in Add WHERE col IS NOT NULL
ORDER BY columns for ORDER BY columns
Result Checker Empty or meaningless Queries returning only NULL
results Zero-row result sets